OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (2, 7) and (-6, 0).

jimthompson5910 (jim_thompson5910):

slope: m = (y2 - y1)/(x2 - x1) m = (0 - 7)/(-6 - 2) m = -7/(-8) m = 7/8

jimthompson5910 (jim_thompson5910):

Now use this slope, and the first point, to find the y-intercept to find the slope intercept form y = mx + b y = (7/8)x + b 7 = (7/8)(2) + b 7 = 14/8 + b 7 - 14/8 = b 7*(8/8) - 14/8 = b 56/8 - 14/8 = b (56 - 14)/8 = b 42/8 = b 21/4 = b b = 21/4 So the equation in slope intercept form is y = 7/8x + 21/4 with me so far? let me know when you want me to continue (to convert to standard form)

OpenStudy (anonymous):

okay yeah i think i follow

jimthompson5910 (jim_thompson5910):

any parts you're stuck on?

OpenStudy (anonymous):

no i understand slope intercept form well, just not standard form

jimthompson5910 (jim_thompson5910):

alright, let's continue then

jimthompson5910 (jim_thompson5910):

the LCD of 7/8 and 21/4 is 8

jimthompson5910 (jim_thompson5910):

so multiply EVERY term by the LCD 8 to get y = 7/8x + 21/4 8*y = 8*(7/8x) + 8*21/4 8y = 56/8x + 168/4 8y = 7x + 42 now move over the x term 8y = 7x + 42 8y - 7x = 42 -7x + 8y = 42 finally, make the x term positive and you do this by multiplying both sides by -1 -7x + 8y = 42 -1*(-7x + 8y) = -1*42 7x - 8y = -42

jimthompson5910 (jim_thompson5910):

so the equation in standard form is 7x - 8y = -42

OpenStudy (anonymous):

thats not an answer option

jimthompson5910 (jim_thompson5910):

hmm what are your choices?

OpenStudy (anonymous):

-7x + 8y = 42 -7x - 8y = 42 -7x + 8y = -42 8x + 7y = 42

jimthompson5910 (jim_thompson5910):

I see

jimthompson5910 (jim_thompson5910):

so they left x alone

jimthompson5910 (jim_thompson5910):

so ignore the last 3 steps

jimthompson5910 (jim_thompson5910):

so what is the final answer

OpenStudy (anonymous):

-7x + 8y = 42?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thank you so much for your help

jimthompson5910 (jim_thompson5910):

you're welcome